OpenStudy (anonymous):

Find y-intersection of x^2-12x+39 I've gotten myself lost -.-

OpenStudy (amistre64):

what is it when x=0?

OpenStudy (anonymous):

oh replace x with 0 to find y-intersection am i right?

OpenStudy (amistre64):

yep; the y axis IS the x=0 axis; so when x=0 ..... where does it hit?

OpenStudy (anonymous):

39

OpenStudy (amistre64):

tada!!

OpenStudy (anonymous):

:D im not very good at this, how about x?

OpenStudy (amistre64):

the x intercepts are when y=0; which is tricker on quadratics

OpenStudy (amistre64):

you gotta have an eye for factors of just plug in the quadratic formula

OpenStudy (amistre64):

x^2-12x+39 = 0 x^2-12x = -39 x^2-12x+36 = -39+36 (x-6)^2 = -3 it dont hit the x axis by this

OpenStudy (anonymous):

ok let me check over thios

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x^2-12x%2B39

OpenStudy (anonymous):

you're awesome!

OpenStudy (amistre64):

.. yep

OpenStudy (anonymous):

Oh where did you get the 36 in the 3rd line above?

OpenStudy (amistre64):

its called completing the square: in a nutshell its just taking that number on the "x" and halving it; then squareing it

OpenStudy (amistre64):